OpenStudy (anonymous):

Help Factor and solve the equation. 9x2 - 30x + 25 = 0 A. {5/3} B. -1/3 C. -3 D. -1

Directrix (directrix):

A. {5/3} 9x^2 - 30x + 25 = 0 (3x -5)(3x-5) = 0 x = 5/3 or x = 5/3 --> double root